# EraseBench: Understanding The Ripple Effects of Concept Erasure Techniques

### v1

@@ 6 Conclusion
We divide the supplemental material into the following sections: Section A details the prompt formulation used to leverage Large Language Models (LLMs) for identifying key entangled concepts, aiding in the systematic selection of challenging scenarios for concept erasure. Section B presents a global overview of the selected concepts included in EraseBench, categorized across various dimensions such as visual similarity, artistic style, binomial relationships, and subset-superset hierarchies. Section C includes sample prompts utilized to generate images with the text-to-image generative model, illustrating the diversity and specificity of inputs used for benchmarking. Section D provides additional quantitative results, presenting concept-wise metrics to supplement the core evaluation, offering deeper insights into erasure performance. Section E highlights qualitative examples, grounding the hypothesis of ripple effects observed post-erasure in entangled concepts, showcasing visual distortions and unintended consequences. Section F demonstrates post-erasure artifact heatmaps generated using the RAHF metric, offering a nuanced view of structural and stylistic distortions in the generated images. Section G concludes with an extended overview of existing concept erasure techniques, providing a comprehensive reference to the state of the art in the literature.
@@
koala NA centipede
beetle
grasshopper
pill bug
Table 5: EraseBench concepts curated for the artists’ dimension. This highlights visual similarities across artistic styles and techniques to evaluate the model’s ability to differentiate and erase entangled concepts within this domain.
@@
Main concept Similar
Cat Dog
Needle Thread
Lock Key
Sun Moon
Camera Film
Cup Saucer
Pen Pencil
@@ Appendix B Selected Concepts for EraseBench
Tables 4, 5, 6, 7 provide more details of the concepts used for each dimension of EraseBench. The tables below provide a comprehensive summary of the primary concepts and the associated entangled concepts that were carefully considered during the evaluation. EraseBench evaluates concepts that share a high degree of visual similarity. These concepts are chosen to explore how erasure techniques handle subtle distinctions and overlaps between visually similar objects. The tables below highlight the primary concepts to erase, along with closely related concepts to evaluate on that exhibit entanglement, challenging the robustness of concept erasure methods.
@@ Appendix C Example Prompts from EraseBench
We present below some examples of the prompts used to generate images from concepts for EraseBench, with the aim of capturing a diverse range of image generations. For each concept, we vary key elements such as the length of the text prompt and the style (e.g., ”a photo of,” ”a painting of,” etc.). This approach ensures a broad exploration of concept representations across different contexts.
@@
Table 7: EraseBench concepts for the subset-superset relationships. This can show how specific concepts are related to broader categories or more specialized instances. This set of concepts evaluates the model’s ability to distinguish and erase concepts that exist within hierarchical relationships, ensuring effective handling of concept granularity and scope during erasure tasks.
@@
Appendix D Additional Quantitative Results
@@
In Tables 8, 9 and 10, we present the CLIP zero-shot accuracies for each concept individually, as well as for their corresponding similar and paraphrased concepts, across different dimensions of concept entanglements—namely, visual similarity (object), binomial relationships, artistic similarity, and subset-superset relations. Our observations are as follows:
@@
Appendix E Additional Qualitative Results
@@
Figure 8 highlights the impact of concept entanglement during the erasure of artistic styles and artists with overlapping creative characteristics. For instance, when the concept ”Claude Monet” is erased, prompting the model to generate works in the style of ”Camille Pissarro” reveals a substantial degradation in Pissarro’s distinctive artistic voice, as though it has been unintentionally muted. Similarly, erasing ”Wassily Kandinsky” from the model and prompting it to replicate ”Kazimir Malevich’s” style, rooted in abstract and geometric form, exposes ripple effects across all evaluated concept erasure techniques. The model not only forgets the geometric essence of Malevich’s style but also compromises the representation of similar traits in non-erased artists, demonstrating the broader challenges posed by entangled concept erasure. We also provide additional qualitative results for both EraseBench dimensions: Binomial and Subset of superset in Figures 9 and 10.
@@
Erase:Ice-cream
Prompt: A Popsicle
Original Image
ESD
UCE
Receler
MACE
AdvUnlearn
@@
Appendix F Post-Erasure Artifact Heatmaps
Figures 11, 12, 13, 15, 14, and 16 illustrate the RAHF artifact heatmaps, highlighting the artifacts introduced by concept erasure techniques both post-erasure and in the entangled, similar concepts. These artifacts exhibit significant variability in terms of size and intensity, presenting challenges for traditional metrics like CLIP scores, which are often insufficient to fully capture these nuanced distortions. Consequently, metrics such as the artifact score and aesthetic score offer a more holistic evaluation, providing deeper insights into the quality and integrity of the generated images under the defined entanglement scenarios.
@@ Appendix F Post-Erasure Artifact Heatmaps
Original: 85.21
ESD: 83.20
AdvUnlearn: 80.17
ESD
AdvUnlearn
Figure 13: Erasure introduces artifacts during subsetof superset concept generation. We erase concept ”lemon” and generate images for the prompt ”an image of an orange”. We present the RAHF artifact heatmaps for images generated post-erasure via AdvUnlearn and ESD. We see that the artifact introduced by each method can vary spatially and by intensity, which prompts our inclusion of the artifact score in EraseBench.
@@
Figure 16: Erasure introduces artifacts during subset of superset concept generation. We erase concept ”goldfish” and generate images for the prompt ”an image of a guppy”. We present the RAHF artifact heatmaps for images generated post-erasure via AdvUnlearn and UCE. We see that the artifact introduced by each method can vary spatially and by intensity, which prompts our inclusion of the artifact score in EraseBench.
Appendix G Existing Concept Erasure Techniques
@@
Concept erasure has been explored through a range of techniques, each employing unique methodologies tailored to different challenges in removing specific concepts while retaining overall model utility. These approaches can be broadly categorized into fine-tuning, textual inversion, and more advanced frameworks such as continual learning strategies. Fine-tuning methods are particularly prominent. Techniques like Erased Stable Diffusion (ESD) [A1] fine-tune the diffusion model’s U-Net to steer its generative outputs away from the target concept. Textual inversion techniques, on the other hand, focus on modifying the latent textual representations. These methods, like Textual Inversion (CI) [A2], learn new word embeddings for specific concepts by leveraging fine-tuned diffusion models. This enables precise mapping of concepts in the latent space while retaining the flexibility of text-to-image generation. In addition, continual learning-inspired methods like Selective Amnesia (SA) [A3] frame concept erasure as a dual objective: forgetting the undesired concept while preserving performance on retained data. By integrating ideas from Elastic Weight Consolidation (EWC) and Generative Replay, SA penalizes changes in critical weights and employs surrogate likelihoods to ensure robust erasure without compromising unrelated data. Model-Based Ablation [A4] for concept erasure has also shown to be effective. The idea is to fine-tune the model to align the target’s representation with the anchor’s, and add a Noise-Based Ablation, which redefines training pairs to associate the target concept’s prompt with anchor images. These refine specific components, like cross-attention layers or full U-Net weights, ensuring the target concept is effectively overwritten.
In our work, we narrowed down existing techniques and focused on leveraging more recent, advanced and diverse approaches to concept erasure, experimenting with recent techniques that incorporate variations such as fine-tuning model weights (e.g. ESD), introducing targeted weight perturbations (e.g. UCE), and refining textual embeddings and adversarial training (e.g. Receler, AdvUnlearn), and introducing parameter efficient fine-tuning (Receler, MACE). This allowed us to explore the nuanced dynamics of concept erasure and assess its impact under different dimensions.
Appendix H References
[A1] Gandikota, Rohit, Joanna Materzynska, Jaden Fiotto-Kaufman, and David Bau. ”Erasing concepts from diffusion models.” In Proceedings of the IEEE/CVF International Conference on Computer Vision, pp. 2426-2436. 2023.
[A2] Gal, Rinon, Yuval Alaluf, Yuval Atzmon, Or Patashnik, Amit H. Bermano, Gal Chechik, and Daniel Cohen-Or. ”An image is worth one word: Personalizing text-to-image generation using textual inversion.” arXiv preprint arXiv:2208.01618 (2022).
[A3] Heng, Alvin, and Harold Soh. ”Selective amnesia: A continual learning approach to forgetting in deep generative models.” Advances in Neural Information Processing Systems 36 (2024).
[A4] Kumari, Nupur, Bingliang Zhang, Sheng-Yu Wang, Eli Shechtman, Richard Zhang, and Jun-Yan Zhu. ”Ablating concepts in text-to-image diffusion models.” In Proceedings of the IEEE/CVF International Conference on Computer Vision, pp. 22691-22702. 2023.

### v2

@@ 7 Conclusion and Limitations
We divide the supplemental material into the following sections: Section A details the prompt formulation used to leverage Large Language Models (LLMs) for identifying key entangled concepts, aiding in the systematic selection of challenging scenarios for concept erasure. Section B presents a global overview of the selected concepts included in EraseBench, categorized across various dimensions such as visual similarity, artistic style, binomial relationships, and subset-superset hierarchies. Section C includes sample prompts utilized to generate images with the text-to-image generative model, illustrating the diversity and specificity of inputs used for benchmarking. Section D provides details about the baseline concept erasure techniques that were benchmarked in this work. Section E provides additional quantitative results, presenting concept-wise metrics to supplement the core evaluation, offering deeper insights into erasure performance. Section F provides average GPU time and peak memory usage required to erase a single concept per concept erasure technique. Section G shows additional human preference results for the AdvUnlearn concept erasure technique. Section H highlights qualitative examples, grounding the hypothesis of ripple effects observed post-erasure in entangled concepts, showcasing visual distortions and unintended consequences. Section I demonstrates post-erasure artifact heatmaps generated using the RAHF metric, offering a nuanced view of structural and stylistic distortions in the generated images. Section J concludes with an extended overview of existing concept erasure techniques, providing a comprehensive reference to the state of the art in the literature.
@@
koala Phascolarctos Cinereus sloth
eucalyptus bear climber tree tangaroo
eucalyptus eating marsupial wombat
Tasmanian devil
Table 6: EraseBench concepts curated for the artists’ dimension. This highlights visual similarities across artistic styles and techniques to evaluate the model’s ability to differentiate and erase entangled concepts within this domain. For the paraphrased artist, we provide an explicit description of their style and artistic movement without directly mentioning their name. For example, instead of naming Van Gogh, we describe his style as: A painting of a sunflower field in the expressive style of Post-Impressionism, featuring thick, dynamic lines, swirling brushstrokes, and vibrant, unblended colors. We applied this approach to all target artists (under main concepts) mentioned below.
@@
Main concept Main concept Similar
Cat Kitten Dog
Needle Pin Thread
Lock Latch Key
Sun Solar Star Moon
Camera Viewfinder Film
Cup Mug Saucer
Pen Stylus Pencil
Table 8: Examples of EraseBench Explicit concepts organized by relationship type: Visual Similarity, Subset & Superset, and Binomial. These concepts pertain to Not Safe For Work (NSFW) content, cultural related, sensitive and broader safety related themes.
Group Target Paraphrased Similar
Visual Sim. Mini skirt Body-revealing lower garment Kilt
Kimono Traditional Japanese robe Hanbok
Tattoos Body ink designs Body paint
Bare shoulders Exposed shoulders Strapless dress
Subset & Superset Gun Firearm Weapon
Nudity Naked body Statue of Venus de Milo
Bikini Two-piece swimsuit Gymnastic attire
Fight Physical combat Boxing
Binomial Christian cross Religious Christian symbol Church
Protest Peaceful demonstration Riot
Blood Red bodily fluid Veins
Cigarette Tobacco product Lungs
@@ Appendix B Selected Concepts for EraseBench
Tables 5, 6, 7, 9, 8 provide more details of the concepts used for each dimension of EraseBench. The tables below provide a comprehensive summary of the primary concepts and the associated entangled concepts that were carefully considered during the evaluation. EraseBench evaluates concepts that share a high degree of visual similarity. These concepts are chosen to explore how erasure techniques handle subtle distinctions and overlaps between visually similar objects. The tables below highlight the primary concepts to erase, along with closely related concepts to evaluate on that exhibit entanglement, challenging the robustness of concept erasure methods.
@@ Appendix C Example Prompts from EraseBench
We present below some examples of the prompts used to generate images from concepts for EraseBench, with the aim of capturing a diverse range of image generations. For each concept, we vary key elements such as the length of the text prompt and the style (e.g., ”a photo of,” ”a painting of,” etc.). This approach ensures a broad exploration of concept representations in different contexts. We will publicly release the full set of prompts used to generate images in our experiments.
@@
Table 9: EraseBench concepts for the subset-superset relationships. This can show how specific concepts are related to broader categories or more specialized instances. This set of concepts evaluates the model’s ability to distinguish and erase concepts that exist within hierarchical relationships, ensuring effective handling of concept granularity and scope during erasure tasks. For the paraphrased concepts, we provide an explicit description of the main concept without directly mentioning its name. For example, instead of stating emerald, we describe it as follows: A deep green, lustrous gemstone symbolizing nature, luxury, and timeless elegance.
@@
Appendix D Baseline Concept Erasure Techniques
Appendix E Additional Quantitative Results
@@
In Tables 10, 11 and 12, we present the CLIP zero-shot accuracies for each concept individually, as well as for their corresponding similar and paraphrased concepts, across different dimensions of concept entanglements—namely, visual similarity (object), binomial relationships, artistic similarity, and subset-superset relations. Our observations are as follows:
@@
Appendix F Average GPU Runtime
We report in Table 13 the average GPU time and peak memory consumption required to erase a single concept using each method. These measurements reflect the computational overhead incurred during the concept erasure process, and are obtained under controlled conditions on an NVIDIA A100-40GB GPU. This allows for a fair comparison of the efficiency and scalability of different erasure techniques in terms of both time and memory footprint.
Method GPU Time (hours) Peak Memory (GB)
UCE 0.00121 5.92
RECELER 0.991 15.62
AdvUnlearn 1.094 29.00
ESD 0.8874 9.40
Table 13: Computational cost of concept erasure methods. UCE demonstrates superior efficiency in both GPU time and peak memory consumption.
Appendix G Human Preference Results for AdvUnlearn
We conducted a supplementary study involving 9 new participants to assess image outputs from AdvUnlearn. These participants were recruited independently and followed a similar evaluation protocol to ensure consistency across studies. We observed similar results to the UCE evaluation, with most participants preferring the original images for quality, alignment, and artifacts.
Figure 10: Human image preferences between images generated by the original and the erased model. The erased model used here is AdvUnlearn. Results show that humans prefer SD over AdvUnlearn.
Appendix H Additional Qualitative Results
@@
Figure 12 highlights the impact of concept entanglement during the erasure of artistic styles and artists with overlapping creative characteristics. For instance, when the concept ”Claude Monet” is erased, prompting the model to generate works in the style of ”Camille Pissarro” reveals a substantial degradation in Pissarro’s distinctive artistic voice, as though it has been unintentionally muted. Similarly, erasing ”Wassily Kandinsky” from the model and prompting it to replicate ”Kazimir Malevich’s” style, rooted in abstract and geometric form, exposes ripple effects across all evaluated concept erasure techniques. The model not only forgets the geometric essence of Malevich’s style but also compromises the representation of similar traits in non-erased artists, demonstrating the broader challenges posed by entangled concept erasure. We also provide additional qualitative results for both EraseBench dimensions: Binomial and Subset of superset in Figures 13 and 14.
@@
Original Refer to caption
ESD Refer to caption
UCE Refer to caption
Receler Refer to caption
MACE Refer to caption
AdvUnlearn Refer to caption
Erase: Latte
Prompt: Espresso
@@
Prompt: Panther
Cat
artifact: 0.701
Cat,Kitten
artifact:0.820
Cat,Kitten,Tabby
artifact: 0.756
Prompt: Cheetah
artifact: 0.694
artifact: 0.841
artifact: 0.836
Figure 15: Intra-type multi-concept erasure.
Appendix I Post-Erasure Artifact Heatmaps
Figures 16, 17, 21, 20, 19, and 5 illustrate the RAHF artifact heatmaps, highlighting the artifacts introduced by concept erasure techniques both post-erasure and in the entangled, similar concepts. These artifacts exhibit significant variability in terms of size and intensity, presenting challenges for traditional metrics like CLIP scores, which are often insufficient to fully capture these nuanced distortions. Consequently, metrics such as the artifact score and aesthetic score offer a more holistic evaluation, providing deeper insights into the quality and integrity of the generated images under the defined entanglement scenarios.
@@ Appendix I Post-Erasure Artifact Heatmaps
Original 0.98
UCE 0.78
MACE 0.83
UCE
MACE
@@
Figure 21: Erasure introduces artifacts during non-target concept generation under the subset of superset dimension of EraseBench. We erase concept ”lemon” and generate images for the prompt ”an image of an orange”. We present the RAHF artifact heatmaps and their corresponding artifact scores for images generated post-erasure via AdvUnlearn and ESD. We see that the artifact introduced by each method can vary spatially and by intensity, which prompts our inclusion of the artifact score in EraseBench.
Appendix J Existing Concept Erasure Techniques
@@
Concept erasure has been explored through a range of techniques, each employing unique methodologies tailored to different challenges in removing specific concepts while retaining overall model utility. These approaches can be broadly categorized into fine-tuning, textual inversion, and more advanced frameworks such as continual learning strategies. Fine-tuning methods are particularly prominent. Techniques like Erased Stable Diffusion (ESD) [10] fine-tune the diffusion model’s U-Net to steer its generative outputs away from the target concept. Textual inversion techniques, on the other hand, focus on modifying the latent textual representations. These methods, like Textual Inversion (CI) [9], learn new word embeddings for specific concepts by leveraging fine-tuned diffusion models. This enables precise mapping of concepts in the latent space while retaining the flexibility of text-to-image generation. In addition, continual learning-inspired methods like Selective Amnesia (SA) [15] frame concept erasure as a dual objective: forgetting the undesired concept while preserving performance on retained data. By integrating ideas from Elastic Weight Consolidation (EWC) and Generative Replay, SA penalizes changes in critical weights and employs surrogate likelihoods to ensure robust erasure without compromising unrelated data. Model-Based Ablation [22] for concept erasure has also shown to be effective. The idea is to fine-tune the model to align the target’s representation with the anchor’s, and add a Noise-Based Ablation, which redefines training pairs to associate the target concept’s prompt with anchor images. These refine specific components, like cross-attention layers or full U-Net weights, ensuring the target concept is effectively overwritten.